# Filtrated Grouping in Multiple Functional Regression

### v2

@@ 4.2 Forest Structure Identification
We set $N=400$ and $\sigma=0.5$. The identified forest structure is given in Table 2. The associated filt-fPLS basis functions are displayed in Figure 5. The filt-fPLS basis functions reveal the common kinematics patterns shared by joints in the same group. For instance, global homogeneity is positively driven by the collective influence of covariates during the intervals 0–0.4 and 0.95–1 of the domain. It is noted that only in the first layer all covariates are clustered into the same group. This global homogeneous component is driven by the first two layers of the oracle structure. In the second layer, the 4th to 10th covariates are clustered together due to their closer values of $b_{jd}$. Specifically, as $d=3,4$, $b_{jd}=1.5$ for $j=4,\ldots,8$ and $b_{jd}=1.8$ for $j=9,10$, and these values are relatively closer to each other compared to $b_{jd}=1$ for $j=1,2,3$. In the 3rd and 4th layers, the closeness between $\beta_{4}(t)$ and $\beta_{5}(t)$, $\beta_{7}(t)$ and $\beta_{8}(t)$, as well as $\beta_{9}(t)$ and $\beta_{10}(t)$, leads to these pairs of covariates being clustered in the same group. In the final layer, the structure captures the individual contributions of each covariate. Overall, the identified forest structure closely aligns with the oracle forest structure and captures the underlying homogeneity more efficiently, as it comprises only six layers and delivers substantially better predictive performance (see Section 4.3).
@@ 4.3 Model Performance and Comparison
To illustrate the superiority of the new filtrated modeling procedure, we compare the prediction performance of the following methods: 1). MFR based on the identified forest structure (Identified); 2). MFR based on the oracle forest structure (Oracle); 3). Grouped multiple function regression (Grouped, see 2-3); 4).  MFR based on the naive grouping structure (NPLS), where no covariates are clustered together in all layers. We test this naive structure to show the necessity of grouping; 5). Ordinary multiple functional linear regression (MFLR).
We repeat 300 simulation runs for each simulation setup. In each run, we generate either 100 or 300 samples for the training set to train the model, and 200 additional samples in the test set used for prediction. The prediction mean squared error (MSE), calculated as $\text{MSE}=\sum_{n=1}^{200}(\hat{y}_{n}-y_{n})^{2}/200,$ is computed for each simulation run. As a result, a total of 300 prediction MSE values are obtained.
The box plots of prediction MSEs are presented in Figure 6. They clearly demonstrate that the identified grouping structure outperforms the oracle structure in capturing the relationship between covariates and the response. Its substantial superiority over the naive structure underscores the importance of grouping. As a baseline, the ordinary multiple functional regression model performs the worst. Although the grouped model performs reasonably well, it still falls short compared to the filtrated grouped model. This is because the homogeneity among functional covariates is multi-resolutional, making a fixed grouping structure across different layers inappropriate. Interestingly, although the data are generated based on the oracle structure, the identified forest structure yields significantly better performance. This improvement arises because the data-driven structure more effectively captures the underlying mechanisms of the data. In particular, covariates exhibiting subtle nuances may be better modeled within the same group, thereby rendering the oracle structure suboptimal.
@@
5 Effective Age and Angular Kinematics
In this study, we explore the relationship between the chronological age and the angular kinematics for the healthy participants in the experiment (see Fukuchi et al. [15]), driven by its relevance to communicating chronic risk analysis. The aim of this study is to develop a model-based system for evaluating the “effective age” of joints, which reflects joint functional health and aging.
5.1 Data Summary, Pre-processing, and Modeling Setup
We average the angular kinematic functions collected from the left and right sides to reduce model complexity. Consequently there are 15 functional covariates, including pelvic obliquity, pelvic rotation, pelvic tilt, hip add/abduction, hip int/external rotation, hip flexion/extension, knee add/abduction, knee int/external rotation, knee flx/extension, ankle inv/eversion, ankle add/abduction, ankle dorsi/plantarflexion, foot inv/eversion, foot int/external rotation and foot DF/plantarflexion angles (see Figure 7). To facilitate the forest structure identification and improve model interpretability, we centralize the log-transformed age and use it as the response, and we scale the covariates (angular kinematics curves) as follows,
$\widetilde{X}_{jn}(t)=\frac{X_{jn}(t)-\overline{X}_{j}(t)}{\sqrt{{N^{-1}}\sum\limits_{n=1}^{N}\|X_{jn}(t)-\overline{X}_{j}(t)\|^{2}}},$
where $\overline{X}_{j}(t)=N^{-1}\sum_{n=1}^{N}X_{jn}(t)$. The participants were required to walk on treadmill under different gait speeds, which are categorized into four groups: 40–55% (S1), 70–85% (S2), 100–115% (S3), and 130–145% (S4) of the normal walking speed. Then we establish models under different gait speeds. Due to page limitations, the results, except those in Section 5.4, focus on the 100–115% normal walking speed range.
Figure 7: Joint angular kinematics functions averaged across the left and right sides, with curve colors indicating walking speed as shown in Figure 1.
5.2 Forest Structure Identification
We employ a Monte Carlo cross-validation approach to select the tuning parameters in the GIC criterion. Specifically, the data is randomly split into a training set (80%) and a test set (20%). The model trained on the training set is used to predict the responses in the test set. We repeat this procedure 1000 times to obtain the average prediction MSE. The tuning parameters $\theta$ associated with the minimal prediction MSE is then used for forest structure identification, resulting in the forest structure displayed in Figure 8. The group indices are displayed as follows:
  • ~$\widetilde{\mathcal{F}}_{1}$.
@@ 5.2 Forest Structure Identification
The forest structure reveals that the associations of pelvis obliquity and hip adduction/abduction with age differ substantially from those of other joints, as they are separated at the first layer. This separation reflects their unique adaptations to age-related changes in gait and balance control. In addition, knee add/abduction also demonstrates significant distinctiveness, leading to its separation in the second layer. In contrast, the kinematics of most other joints exhibit a high degree of interdependence, suggesting more uniform adaptations to preserve forward propulsion and overall gait efficiency.
5.3 Homogeneous Components
The filt-fPLS basis functions $\psi_{d,i}(t)$ reveal how the homogeneous components are associated with age. Figure 9 presents the filt-fPLS basis functions in the first two layers. These basis functions reflect how the homogeneous components are affected by the relevant angular kinematics. For instance, in the first layer, the homogeneous component of group 1 is positively determined by the related angular kinematics (pelvis obliquity and hip adduction/abduction) during 40–70% of the gait cycle and negatively affected during the remainder of the gait cycle. The homogeneous components of group 2 (driven by all other joints) is uniformly positively determined by the related angular kinematics.
@@ 5.3 Homogeneous Components
Figure 9: The partial least square basis functions in the first two layers $\{\psi_{d,i}(t)\colon d=1,2,\ i\geq 1\}$.
Define the model component of the $i$-th group in the $d$-th layer as $a_{d,i}\zeta^{(i)}_{nd}$. We display the 95% confidence interval of the coefficient scores $\{a_{d,i}\colon i,d\geq 1\}$ and the standard deviation of the associated model components in Figure 10. The upper panel, which shows the confidence interval of the coefficient scores, highlights how different homogeneous components are associated with the response. The homogeneous components in the first layer, corresponding to the two sets of joints, both exhibit positive associations with the response, and in the second layer, all homogeneous components except for the one related to hip add/abduction are also positively associated with the response.
Interestingly, the results reveal a prominent spike in variation for hip flexion/extension. This suggests that the idiosyncratic gait patterns of hip flexion/extension plays a crucial role in reflecting the effective age. As people grow older, there’s a well-documented distal‑to‑proximal shift in gait: propulsion and control move away from ankle and knee toward hip. This shift makes hip movement a central driver of gait, so its variability becomes especially telling (see e.g., Franz and Kram [14]). Certain homogeneous components are found to be non-significant in the model. Specifically, the idiosyncratic components in the third layer associated with pelvic obliquity, pelvic rotation, knee adduction/abduction, knee internal/external rotation, knee flexion/extension, and ankle dorsiflexion/plantarflexion show negligible associations with age.
[…]
5.4 Predictive Power
In this section, we evaluate the predictive performance of the filtrated model, developed under various gait speed levels categorized in Section 5.1, in predicting the chronological age. We compare the proposed method with two competing approaches: the grouped functional regression model (2-3) and the ordinary functional regression model (1-1). We use the same Monte Carlo cross-validation approach to evaluate predictive performance, and the prediction MSEs are shown in Figure 11. The results indicate that the filtrated model achieves the best performance at 70–80% of normal walking speed, suggesting that participants should be encouraged to walk at a slightly slower pace than their natural, comfortable speed when assessing their joint effective age.
@@ 5.4 Predictive Power
Figure 11: Box-plots of MSEs of different methods. The variance of the response variable under the four gait speed levels is 0.19, 0.19, 0.19, and 0.17, respectively.
The filtrated model consistently provides the most accurate and robust predictions. In contrast, the ordinary model underperforms due to its high complexity, which leads to substantial estimation errors and diminished predictive power. While the grouped model demonstrates improved performance over the ordinary model, it remains suboptimal compared to the filtrated model. This discrepancy arises because the idiosyncratic components across different covariates exhibit significant variation. Treating these heterogeneous components collectively, as in the grouped model, fails to capture their distinct contributions, thereby limiting predictive accuracy.
6 Conclusion
@@
We propose a novel filtrated grouping procedure for multiple functional regression models with a scalar response and multiple functional covariates, addressing the challenges posed by the joint coordination and joint effective age evaluation. By introducing a forest-structured grouped model and a data-driven filtrated partial least squares estimation procedure, our approach effectively captures both homogeneous and idiosyncratic effects of covariates from a multi-resolution perspective. Our results demonstrate the potential of this framework to uncover biomechanical adaptations and mechanisms associated with joint aging. By hierarchically grouping covariates, the method has the ability to identify both homogeneous and idiosyncratic features across different joints provides a comprehensive understanding of the interdependence in gait dynamics, supporting the evaluation of effective age and facilitating informed decision-making in clinical and research settings. The proposed approach surpasses existing methods in both interpretability and predictive accuracy, and provides a comprehensive modeling framework for complex covariate relationships.
This study marks the first effort to integrate filtration techniques into multiple functional regression, and our results underscore the substantial potential and impact of this approach. The framework is inherently generalizable and can be extended to other applications involving multiple functional covariates. Future work will broaden the filtration paradigm to other functional methodologies—such as functional factor models and classification frameworks.

### v3

@@ 3.2 Forest Structure Identification
We set $N=400$ and $\sigma=0.5$. The identified forest structure is given in Table 2. The associated filt-fPLS basis functions are displayed in Figure 5, revealing the common variation patterns which determine the associated homogeneous components. For instance, global homogeneity is positively driven by the collective influence of covariates during the intervals 0–0.4 and 0.95–1 of the domain. It is noted that only in the first layer all covariates are clustered into the same group. This global homogeneous component is driven by the first two layers of the oracle structure. In the second layer, the 4th to 10th covariates are clustered together due to their closer values of $b_{jd}$. Specifically, as $d=3,4$, $b_{jd}=1.5$ for $j=4,\ldots,8$ and $b_{jd}=1.8$ for $j=9,10$, and these values are relatively closer to each other compared to $b_{jd}=1$ for $j=1,2,3$. In the 3rd and 4th layers, the closeness between $\beta_{4}(t)$ and $\beta_{5}(t)$, $\beta_{7}(t)$ and $\beta_{8}(t)$, as well as $\beta_{9}(t)$ and $\beta_{10}(t)$, leads to these pairs of covariates being clustered in the same group. In the final layer, the grouping structure captures the individual contributions of each covariate. Overall, the identified forest structure closely aligns with the oracle forest structure and captures the underlying homogeneity more efficiently, as it comprises only six layers and delivers substantially better predictive performance (see Section 3.3).
@@ 3.3 Model Performance and Comparison
To illustrate the superiority of the new filtrated modeling procedure, we compare the prediction performance of the following methods: 1). MFR based on the identified forest structure (Identified); 2). MFR based on the oracle forest structure (Oracle); 3). Grouped multiple function regression (Grouped); 4).  MFR based on the naive grouping structure (NPLS), where no covariates are clustered together in all layers. We test this naive structure to show the necessity of grouping; 5). Ordinary multiple functional linear regression (MFLR).
We repeat 300 simulation runs for each simulation setup. In each run, we generate either 100 or 300 samples for the training set to train the model, and 200 additional samples in the test set used for prediction. The prediction mean squared error (MSE), calculated as $\text{MSE}=\sum_{n=1}^{200}(\hat{y}_{n}-y_{n})^{2}/200,$ is computed for each simulation run. As a result, a total of 300 prediction MSE values are obtained. The box plots of prediction MSEs are presented in Figure 6. They clearly demonstrate that the identified grouping structure outperforms the oracle structure in capturing the relationship between covariates and the response. Its substantial superiority over the naive structure underscores the importance of grouping. As a baseline, the ordinary multiple functional regression model performs the worst. Although the grouped model performs reasonably well, it still falls short compared to the filtrated grouped model. This is because the homogeneity among functional covariates is multi-resolutional, making a fixed grouping structure across different layers inappropriate. Interestingly, although the data are generated based on the oracle structure, the identified forest structure yields significantly better performance. This improvement arises because the data-driven structure more effectively captures the underlying mechanisms of the data. In particular, covariates exhibiting subtle nuances may be better modeled within the same group, thereby rendering the oracle structure suboptimal.
@@
4 Filtrated Kinematic Connectivity Analysis
Motivated by its relevance to communicating chronic lower-limb risk associated with aging, we apply the proposed filtrated grouping framework to investigate the relationship between chronological age and angular kinematics using the data collected from the experiment conducted by Fukuchi et al. [14]. Because the sample size is limited, we average the angular kinematic trajectories from the left and right sides. This reduces model complexity while retaining focus on kinematic connectivity. Consequently, there are 15 kinematic trajectories per gait cycle, each serving as a functional covariate, including pelvic obliquity, pelvic rotation, pelvic tilt, hip add/abduction, hip int/external rotation, hip flexion/extension, knee add/abduction, knee int/external rotation, knee flx/extension, ankle inv/eversion, ankle add/abduction, ankle dorsi/plantarflexion, foot inv/eversion, foot int/external rotation and foot DF/plantarflexion angles (see Figure 7). To improve model interpretability, we centralize the log-transformed age and use it as the response $y_{n}$, and standardize the angular kinematic curves as follows,
$\widetilde{X}_{jn}(t)=\frac{X_{jn}(t)-\overline{X}_{j}(t)}{\sqrt{{N^{-1}}\sum\limits_{n=1}^{N}\|X_{jn}(t)-\overline{X}_{j}(t)\|^{2}}},\ j=1,\ldots,15,$
where $\overline{X}_{j}(t)=N^{-1}\sum_{n=1}^{N}X_{jn}(t)$. This preprocessing ensures that group formation reflects the homogeneity of structural and temporal movement patterns, rather than being biased by differences in numerical angular range.
Figure 7: Angular kinematic trajectories averaged across the left and right sides, with curve colors indicating walking speed as shown in Figure 1.
Here, we focus on the comfortable walking speed, specifically within the range of 100–115% of normal walking speed. We employ a Monte Carlo cross-validation approach (see Xu and Liang [37]) to select the tuning parameters in the GIC criterion. Specifically, the samples are randomly split into a training set (80%) and a test set (20%). The model trained on the training set is used to predict the responses in the test set. We repeat this procedure 1000 times to obtain the average prediction MSE. The tuning parameters $\theta$ associated with the minimal prediction MSE is then used for forest structure identification, resulting in the forest structure displayed in Figure 8. The group indices across different filtration layers are shown below:
  • $\widetilde{F}_{1}$.
@@ 4 Filtrated Kinematic Connectivity Analysis
Define the partial sum of squares (PSS) of the homogeneous components of layer $d$ group $i$ as $\mbox{RSS}_{d,i}$, where $\mbox{RSS}_{d,i}$ represents the residual sum of squares of the reduced model with the corresponding homogeneous component omitted from the full model (all the identified homogeneous components are involved). We use the partial sum of squares to evaluate the importance of the homogeneous components, and a larger value of PSS indicates greater importance. We display the 95% confidence interval of the coefficient scores $\{a_{d,i}\colon i,d\geq 1\}$ and the partial sum of squares of all the identified homogeneous components in Figure 9. It should be noted that all coefficient scores in the first two layers are significantly non‑zero, confirming our hypothesis that kinematic connectivity imparts a consistent age-associated pattern across different kinematic trajectories.
Figure 9: The confidence intervals for coefficient scores and the PSS values. Different layers are separated by blue dashed lines. LdGi represents the $i$-th group in the $d$-th layer, same as $\mathcal{K}_{d,i}$. The insignificant homogeneous components are marked in red.
The homogeneity of the age-kinematics association is captured by the homogeneous components in $\mathcal{K}_{1,1}$, $\mathcal{K}_{1,2}$, and $\mathcal{K}_{2,2}$. In the first hierarchical layer, covariate 1 and 4, representing pelvic obliquity and hip adduction/abduction in the frontal plane, are separated from the other angular kinematic trajectories. Their associated filt‑fPLS basis function (see Figure 10) exhibit pronounced divergence from the one of the other group in the same layer, underscoring that pelvic and hip frontal-plane dynamics exhibit distinct association with age. This reflects the corresponding collective neuromuscular and postural adaptations: during aging, decreased mediolateral balance control and altered weight distribution drive compensatory adjustments in pelvis and hip frontal-plane motion. The corresponding PSS value is considerably higher than those of other homogeneous components, indicating that pelvic and hip frontal-plane motion coordinate as the primary kinematic feature for assessing joint effective age. Meanwhile, we find that the idiosyncratic kinematic features (pertaining to the homogeneous components identified for groups containing a single covariate) of knee internal/external rotation, knee flexion/extension, and ankle dorsiflexion/plantarflexion are non-significant in their associations with age. Their associations with age are adequately captured in the first two filtration layers through the homogeneous components shared with other kinematic trajectories, and their idiosyncratic components do not distinguish gait characteristics between younger and older individuals. Interestingly, the PSS value of the homogenous components in $\mathcal{K}_{3,11}$, corresponding to the idiosyncratic feature of ankle adduction/abduction dynamics exhibits a pronounced spike. Additionally, the three homogeneous components in $\mathcal{K}_{3,11}$, $\mathcal{K}_{4,4}$, and $\mathcal{K}_{5,1}$, which also correspond to the idiosyncratic features of ankle adduction/abduction, all show significant effects. These findings suggest that ankle adduction/abduction kinematics also serve as a particularly informative biomarker for evaluating effective age, and its individual adaption to aging is non-negligible.
Here we summarize the key findings in the filtrated grouping structure:
  • Different angular kinematic trajectories significantly exhibit homogeneity in their age-related association patterns.
  • Pelvic obliquity and hip adduction/abduction kinematics coalesce into a distinct homogeneous component, coordinately serving as the primary indicator of lower-limb joint aging and revealing the age-associated adaptations in pelvic and hip frontal-plane motions linked to mediolateral balance control.
  • Other kinematic trajectories show more homogeneous adaptive patterns, pointing toward system‑wide biomechanical changes associated with aging in gait. Idiosyncratic features of different angular kinematic trajectories are also pronouncedly associated with aging, except for knee internal/external rotation, knee flexion/extension, and ankle dorsiflexion/plantarflexion.
The filt-fPLS basis functions $\{\psi_{d,i}(t)\colon d\geq 1,\ i=1,\ldots,m_{d}\}$ are displayed in Figure 10, and provide interpretable insights into how the homogeneous components relate to age over the gait cycle. Figures 9 and 10 provide a clear interpretation of how the corresponding kinematic trajectories are associated with age. For example, in the first layer, the homogeneous component in Group 1 is primarily associated with pelvis obliquity and hip adduction/abduction. It exhibits a positive relationship with these angular kinematic trajectories during 40–70% of the gait cycle and a negative relationship during the remaining phases, forming a bell-shaped pattern. Combining the result that the associated coefficient score is significantly positive, we conclude that the 40–70% period of the aggregated kinematic trajectories is positively associated with age, whereas the remaining phases exhibit a negative association. In contrast, the homogeneous component in Group 2 are positively driven by all other angular kinematic trajectories throughout the entire gait cycle and exhibit positive association with the response. The corresponding filt-fPLS basis is approximately a constant function, capturing the mean level of the aggregated kinematic trajectories.
@@
Figure 10: The partial least square basis functions in the first two layers.
We benchmark the performance of the new methods against two established alternatives: the ordinary functional regression model (1-1) and the grouped functional regression model (2-3), through comparing the prediction error for age. Prediction accuracy is evaluated using Monte Carlo cross-validation. The participants were required to walk on treadmill under different gait speeds, and we categorize them into four groups: 40–55% (S1), 70–85% (S2), 100–115% (S3), and 130–145% (S4) of the normal walking speed. Then we establish multiple functional linear models under different gait speed groups with response $y^{(s)}_{n}$ and covariates $\widetilde{X}^{(s)}_{jn}(t)$, where the superscript $(s)$ indicates the gait speed levels, $y^{(s)}_{n}=\sum_{j\geq 1}\langle\widetilde{X}^{(s)}_{jn},\beta^{(s)}_{j}\rangle+\epsilon_{n}^{(s)}.$ We repeat the Monte-Carlo sampling 1000 times, and the resulting mean squared errors (MSEs) are displayed in Figure 11. The filtrated model achieves optimal predictive performance at 70–85% of normal walking speed, indicating that assessments conducted at a slightly reduced pace relative to an individual’s comfortable walking speed yield the most precise effective age assessment.
Across all gait speed levels, the filtrated model consistently delivers the lowest and most robust MSEs. The ordinary model (1-1) underperforms due to its high parameter complexity, which leads to increased assessment uncertainty. The grouped model (2-3) demonstrates modest improvements over the ordinary model; however, it remains inferior to the filtrated model. This shortfall stems from its failure to accommodate heterogeneous, idiosyncratic associations between individual covariates and response. By grouping covariates consistently over all dimensions, the grouped model masks their distinct contributions. In contrast, the filtrated model’s hierarchical grouping structure effectively distinguishes shared versus covariate-specific effects, allowing for much more nuanced modeling of age-kinematics associations and significantly enhancing both interpretability and evaluation performance.
@@
Figure 11: Box-plots of MSEs of different methods. The variance of the log-transformed age under the four gait speed levels is 0.19, 0.19, 0.19, and 0.17, respectively.
5 Discussion
@@
We propose a novel filtrated grouping procedure for multiple functional regression models with a scalar response and multiple functional covariates, addressing the challenges posed by lower-limb joint effective age evaluation. By introducing a forest-structured grouped model and a data-driven filtrated partial least squares estimation procedure, our approach effectively captures both homogeneous and idiosyncratic effects of covariates from a multi-resolution perspective, providing systemic insights into kinematic connectivity in relation to age. The filtrated grouping structures are inferred from data, rather than being imposed a priori, making the grouping structures effectively reflect the data mechanism. We apply the filtrated grouping framework to investigate the relationship between chronological age and joint motion patterns in healthy subjects, and identify several key factors that reflect joint aging mechanisms both collectively and individually across different kinematic trajectories. Our results demonstrate the potential of this framework to uncover biomechanical adaptations associated with joint aging, supporting the evaluation of lower-limb effective age and facilitating informed decision-making in clinical and research settings. The proposed approach surpasses existing methods in both interpretability and predictive accuracy, and provides a stratified modeling framework to account for complex covariate connectivity.
Since the filtrated grouping structure is multi-resolution, we should not focus on any single layer but instead consider the overall architecture of grouping. If two angular kinematic trajectories are highly connected, they tend to cluster together across many filtration layers. This insight is useful because persistent grouping across multiple layers signals truly meaningful biomechanical relationships, rather than artifacts of a single grouping structure. In other words, kinematic trajectories that remain clustered consistently across multiple layers of filtration robustly unravels the strong connections.
There are some limitations of our proposed approach. Since the grouping structure follows a forest structure, and becomes increasingly complicated when the dimension increases, it is difficult to provide comprehensive theoretical analysis. In addition, since coefficient homogeneity is sensitive to covariate scaling, the grouping structure can vary when the covariates are scaled differently. This limitation is inherent in any grouping algorithm that relies on coefficient equality. Although covariate standardization can help stabilize group identification, it does not fully eliminate this sensitivity. In contrast, Jiao and Chan [17] developed a scaling-invariant grouping pursuit algorithm based on coefficient shape homogeneity instead of coefficient equality, termed coefficient shape alignment, in which the grouping structure is invariant to covariate scaling without requiring any data preprocessing. We will integrate our filtration-based analysis with coefficient shape alignment in future work to resolve this issue.
This study marks the first effort to integrate filtration techniques into multiple functional regression, and our results underscore the substantial potential and impact of this approach. This novel analytical framework opens up a number of promising directions for future research. For example: How does multi-resolution kinematic connectivity vary across different gait speeds? Are there significant differences in this connectivity structure between normotypical and pathological gait? By comparing the filtrated grouping structures obtained with the same tuning parameters across different samples, we obtain a more comprehensive and insightful comparison. A similar idea has also been applied in network analysis (see Lee et al. [22]) to obtain stable network comparison across different filtration scales. The framework is inherently generalizable and can be extended to other applications involving multivariate functional data. Future work will broaden the filtration paradigm to other functional data methodologies, such as functional factor models and multivariate functional data classification.